OpenStudy (anonymous):

Find the arc length of the curve y = ln (cos x) for 0 < x < pi/4

OpenStudy (amistre64):

\[\int_{0}^{\pi/4}\sqrt{1+[f'(x)]^2}dx\]

OpenStudy (amistre64):

ln(cos(x))' = -\(\cfrac{sin(x)}{cos(x)}\)

OpenStudy (amistre64):

or simply: -tan(x) \[\int_{0}^{\pi/4}\sqrt{1+[-tan(x)]^2}dx\] \[\int_{0}^{\pi/4}\sqrt{sec^2(x)}dx\] \[\int_{0}^{\pi/4}sec(x)dx\]

OpenStudy (amistre64):

and sec(x) has a naughty little secret ...

OpenStudy (amistre64):

\[ \int_{0}^{\pi/4}sec(x)dx=\int_{0}^{\pi/4}sec(x)\frac{sec(x)+tan(x)}{sec(x)+tan(x)}dx\] \[ \int_{0}^{\pi/4}\frac{sec^2(x)+sec(x)tan(x)}{sec(x)+tan(x)}dx\] \[ \int_{0}^{\pi/4}\frac{sec^2(x)+sec(x)tan(x)}{tan(x)+sec(x)}dx\] now the top is the derivative of the bottom and its just an ln|u| up

OpenStudy (amistre64):

\[ln|tan(\pi/4)+sec(\pi/4)|-ln|tan(0)+sec(0)|\] \[ln|1+\sqrt{2}|-ln|0+1|\] \[ln(1+\sqrt{2})\] maybe?